Perimeter of a triangle is 48 cm. one side has a length of 18 cm. find two other sides of this triangle if their difference is 4.6cm. How do I solve for this?
Answer by macston(5194)

a=one side=18cm; b and c are other sides
b-c=4.6cm
b=4.6cm+c
p=a+b+c
48cm=18cm+b+c Subtract 18cm from each side
30cm=b+c Substitute for b
30cm=(4.6cm+c)+c Subtract 4.6cm from each side
25.4cm=2c Divide each side by 2.
12.7cm=c ANSWER 1: One of the other sides is 12.7cm.
b=4.6cm+c=4.6cm+12.7cm=17.3cm ANSWER 2: The other side is 17.3 cm.
CHECK:
a+b+c=48cm
18cm+12.7cm+17.3cm=48cm
48cm=48cm